Main INDEX
Monthly INDEX
PREV
NEXT
User name jpchen
Log entry time 00:11:35 on October13,2006
Entry number 186955
keyword=swing shift summary
swing shift summary
shift workers: J. P. Chen, V. Tvaskis.
16:00 bema back after arc measurement. position off (1mm, 1mm)
requestred to center it.
16:14, beam cenetered. started production data taking.
18:30-18:44 found right NMR problem. reboot and fixed.
19:03-19:11 found left ioc problem, reboot and fixed.
20:18-20:50 MCC test.
20:45-20:54 Right arm NMR problem, reboot, fixed.
23:00-23:12 Left IOC stuck. Rebooted. Took a few minutes to get dipole to settle.
run 1255 16:14-16:47, 80 uA on He4, 1.5 M events.
run 1256 16:48-17:19, 80 uA, 1.5M.
run 1257 17:20-17:55, 80 uA, 1.5M.
run 1258 17:56-18:28, 80 uA, 1.5M.
run 1259 18:29-18:31, 80 uA, 0.1M, right NMR problem, junk.
run 1269 18:44-19:03, 80 uA, 0.8M, left ioc problem.
run 1261 19:11-19:41, 80 uA, 1.5M.
run 1262 19:42-20:18, 80 uA, 1.5M.
run 1263 20:54-21:25, 80 uA, 1.5M.
run 1264 21:26-21:57, 80 uA, 1.5M.
run 1265 21:58-22:34, 80 uA, 1.5M.
run 1266 22:34-23:00, 80 uA. 1.18M, left ioc stuck near the end of run
run 1267 23:08-23:11, 80 uA, 0.1M, left dipole not settled, junk.
run 1268 23:13-, 80 uA, still on-going at the end of the shift.